Operations Report – April 21st

Crew 192 Operations Report 20Apr2018

SOL 13

Name of person filing report: Richard Blakeman

Non-nominal systems: Upper deck faucet leak continues.

Generator (hours run): 12 hours

Solar ‒ SOC 100%

Diesel – 75% fuel

Propane – 60% (telemetry reported by Shannon) psi.

Ethanol Free Gasoline (5 Gallon containers for ATV) – 25% gas left in one can.

Water (trailer) – Empty water tank secure on Trailer.
Water (static) – 500 gallons (two external tanks)
Trailer to Static Pump used – No
Water (loft) – Static to Loft Pump used – Yes
Water Meter: 133860.0

Toilet tank emptied: Yes. Last emptied today, April 21.

ATV’s Used: None

Oil Added? No

ATV Fuel Used: None

# Hours the ATVs were Used today: 0

Notes on ATVs: All working and ready to use.

Deimos rover used: No (decommissioned)
Hours: 122.3
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Sojourner rover used: ASSIGNED TO DIRECTOR
Hours:
Beginning charge:
Ending charge:
Currently charging:

Spirit rover used: No
Hours: 32.8 (corrected)
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Opportunity rover used: No
Hours: 30.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes

Curiosity rover used: No
Hours: 27.4
Beginning charge: 100%
Ending charge: 100%
Currently charging: No

HabCar used and why, where? Yes, dinner with incoming crew and water runs.

General notes and comments: Last minute cleaning and tasks to prepare hab for incoming crew.

Summary of internet: 0MB at 1900.

Summary of suits and radios: All suits and radios cleaned, charged, and prepped for incoming crew.

Summary of Hab operations: Worked through departure checklist.

Summary of GreenHab operations: See GreenHab Report.

Summary of ScienceDome operations: None

Summary of RAM operations: None

Summary of health and safety issues: None

Questions, concerns, and requests to Mission Support: None